title


NSA OPS OLD NEWS:

  • [02/03/10]: System maintenance
    A major upgrade was performed on the nanuna core operative system. The upgrade went well.

  • [02/27/09]: nsamfrsrC2.a0 and nsamfrsrC2.b1 19990823-20081120 reprocessed
    Please note that the NSA C2 (Atqasuk) MFRSR data for the period 19990823-20081120 has been re-processed. To obtain the new data please go to the ARM Data Management Facility (DMF).

  • [02/02/09]: Replaced faulty power supply
    Many thanks to the Fairbanks "in situ" people, we got the faulty power supply on nanuna replaced. It has caused a series of spontaneous re-boots over the last 3 to 4 months.

  • [07/27/08]: Update on the upgrade
    The upgrade of nanuna is complete now, QL production resumed earlier last week with the arrivial of 64 bit IDL and new license file. We will pass half a million QL images pretty soon too now, but with the new hardware this is easy to handle.

  • [06/26/08]: nanuna has been upgraded
    All brand new hardware, mostly. There are some issues to be worked out because we are now on a 64 bit architechture (previous was 32 bit), and so binaries from old-nanuna does not work on new-nanuna. Expect to have this solved (and QL production running again) by the end of this week.

  • [10/02/06]: Data now from the DMF
    The data we use for Quick-Looks (QLs) and Quality Assurance is now downloaded from the ARM Data Management Facility (DMF). This saves us some bandwidth to the North Slope, but some data get in a bit delayed and in chunks (e.g. the TSI).

  • [09/26/06]: IOP camera from sky-deck
    We have an extra live (near real-time) camera going because of the Radiometer Intercomparison IOP. Our friend the Total Sky Imager (TSI) is also back. Both images can be seen at the Barrow (C1) page. The IOPCam updates every 5 minutes! There is also a daily movie from the TSI images there.

  • [07/24/06]: Disk failure (again)
    Just when we thought we had seen the last of the SCSI disk problems they re-appeared. We removed the entire SCSI sub-system this time, good riddance. Hopefully we will live in peace forever after. No important data was lost but it looked pretty grim for a while.
    We are not back to 100% in terms of backup and redundancy, but we will address this issue soon.
    You might notice temporary outages (e-mail, web, or data-base) as we move things around a bit on the remaining disks.

  • [06/21/06]: Power failure
    One of our UPS systems failed on us and took out both nanuna and agvik. There was some issues with confused disks afterwards, but we worked through it. Web and e-mail services should be back to normal. Thanks again to Dana for being Hans' hands :-)
    Dana will go on vacation starting 6/23, good thing this happened before! Hans is available as usual by e-mail.
    Happy solstice!



  • [03/23/06]: Sync your time with ours
    We have been running a time service (NTP) on nanuna for some time, but we have recently upgraded it and made it available for everyone. You can check on our time service and synchronize your watch here. Accurate time-keeping is vital for collecting data and performing automated tasks.

  • [03/15/06]: Snow-birds are back again
    The multi-filter radiometers (MFRs) have returned to the sites after their winter break in sunnier surroundings. Data are available from about March 15. We are still waitning for news about the TSI.

  • [08/16/05]: Disk failure in redundant array
    We lost a disk in one of our disk arrays. This means our redundancy strategy is working as intended - had this been a lone drive failure, nanuna would be down and I would be on a red-eye to Fairbanks. We are continuing as normal and new disks are under way (thanks Jeff!). Expect minimal down-time as we replace the faulty disk, notices will be sent out. [08/31/05] Update: Today both disks in the disk array were replaced and we are 100% restored. Thanks to Dana!

  • [08/12/05]: Present Weather Sensor (PWS) QL added
    A new Quick Look (QL) has been added for the Present Weather Sensor on the surface meteorological tower (mettwr) in Barrow (C1). Thanks to Dana!

  • [06/23/05]: Nanuna has been upgraded
    Our servers, "nanuna" and its spare twin "anja", has been upgraded. Hopefully you can't notice anything, the upgrade is all behind the scenes stuff. For example, nanuna now has more memory and is about twice as fast after hardware updates. The new systems are based on the Debian "stable" Linux distribution. If you notice any problems, please contact Hans or Dana.

  • [06/20/05]: Another New Addition to the UAF/OPS Team
    We welcome Martin Stuefer into our group!

  • [05/15/05]: New Addition to the UAF/OPS Team
    On Sunday 5/15 at 1:33 am, we got a new member of the NSA/UAF Operations Team. Her name is Sonja, and as she is not much more than 8 lbs 8 oz for the moment, we expect it will be some time before she is ready to wrestle polar bears. Congrats Dana and Martin!

  • [09/16/04]: Firewall installed
    A firewall is now in place on nanuna. If something that used to work no longer does, please contact Hans Eide or Dana Moudry. E-mail and web-services should be entirely unaffected.

  • [06/14/04]: Another disk crash
    After 471(!) days of uninterrupted up-time we had another disk failure on our server nanuna. This time we lost a disk dedicated to backups so we are back and working as normal already.

  • [04/15/04]: New VCEIL and METTWR QLs
    The new met-tower ingest is shaping up and so we have a new stream of quick-looks coming from there. Also check out the new ceilometer quick-looks that are now made using IDL instead of Matlab.

  • [04/09/04]: The return of the *MFRs.
    Our "snowbirds", the MultiFilter Radiometers (MFRs), have returned to the cold north after a prolonged calibration wintering in the south. You can now find quick-looks for these data streams here.

  • [03/11/04]: The NSA OPS web e-mail interface on nanuna has been upgraded to a newer version.
    Please let us know if there are any problems (or if you need an account). Also note that our security certificates has changed.

  • [03/05/04]: There is now a new official NSA Logo that you can download here.
    The new logo reflects that the NSA site is becoming a part of the ACRF Climate Research Sites.

  • [08/11/03]: Level B1 data from the Normal Incidence Multi-channel Radiometer (NIMFR) at both Barrow (C1) and Atqasuk (C2) is now available
    As of 8/10/03, QLs made from only from the new B1 level data. The old QLs based on A0 data are still available of course. Please look for new(old) QLs under the nsanimfrC1.b1(a0) and nsanimfrC2.b1(a0) streams.

  • [05/15/03]: Telephone conference minutes available
    There is now searchable/browsable interface for online NSA OPS telephone conference minutes. We intend to add more documents to this data-base in the future.

  • [04/15/03]: Skyrad and groundrad level B1 data available
    Starting 3/27/03 there are no longer QLs made from a1 data, simply because the ingest is now doing b1 level data and no a1 level data is available anymore. Please look for QLs under the nsaskyrad60sC1.b1 and nsagndrad60sC1.b1 streams.



  • Back to main page